OpenStudy (anonymous):

how do I solve the y=x-9 3y-2x=36 using the substitution method ?

OpenStudy (anonymous):

3(x-9) - 2x = 36 3x - 27 - 2x = 36 3x - 2x = 36 - 27 3x - 2x = 9 x = 9 Is that okay?(: Do you need to find y?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

To find y, sub. the value of x into the equation y = x - 9 x = 9 y = x - 9 y = 9 - 9 y = 0 (:

OpenStudy (anonymous):

thanks, I think I got it

OpenStudy (anonymous):

Okay, you're welcome(:

OpenStudy (anonymous):

I think it's correct.